# TopoStreamer: Temporal Lane Segment Topology Reasoning in Autonomous Driving

### v2

@@ Abstract
Lane segment topology reasoning constructs a comprehensive road network by capturing the topological relationships between lane segments and their semantic types. This enables end-to-end autonomous driving systems to perform road-dependent maneuvers such as turning and lane changing. However, the limitations in consistent positional embedding and temporal multiple attribute learning in existing methods hinder accurate roadnet reconstruction. To address these issues, we propose TopoStreamer, an end-to-end temporal perception model for lane segment topology reasoning. Specifically, TopoStreamer introduces three key improvements: streaming attribute constraints, dynamic lane boundary positional encoding, and lane segment denoising. The streaming attribute constraints enforce temporal consistency in both centerline and boundary coordinates, along with their classifications. Meanwhile, dynamic lane boundary positional encoding enhances the learning of up-to-date positional information within queries, while lane segment denoising helps capture diverse lane segment patterns, ultimately improving model performance. Additionally, we assess the accuracy of existing models using a lane boundary classification metric, which serves as a crucial measure for lane-changing scenarios in autonomous driving. On the OpenLane-V2 dataset, TopoStreamer demonstrates significant improvements over state-of-the-art methods, achieving substantial performance gains of +3.0% mAP in lane segment perception and +1.7% OLS in centerline perception tasks.
Codehttps://github.com/YimingYang23/TopoStreamer
Figure 1: Comparsion between current streaming-based map learning methods (Yuan et al. 2024) and our TopoStreamer. TopoStreamer delivers more fine-grained road information through streaming perception of lane segments, which is vital for planning.
Introduction
@@
Perception serves as a crucial component in end-to-end autonomous driving (Li et al. 2024b; Yang et al. 2025b), providing essential road priors for planning. Existing HD map learning and lane topology reasoning methods primarily focus on frame-by-frame detection (Li et al. 2023b; Liao et al. 2022). This approach fails to account for instance consistency across consecutive frames, making it susceptible to missed detections due to occlusions and high-speed movements (Yuan et al. 2024). Such limitations significantly hinder continuous and smooth decision-making and maneuvers. To comprehensively leverage temporal information, streaming-based methods (Yuan et al. 2024; Wang et al. 2024b; Wu, Yang, and Li 2025) propose memory-based temporal propagation to establish long-term frame associations. Specifically, these approaches leverage the ego-vehicle pose to predict the probable positions of road instances in subsequent frames. However, they fail to capture sufficient road information for planning. This inspired us to introduce a temporal mechanism in lane topology reasoning, which we can leverage perception and topology reasoning results from previous frames to predict current frame outcomes and capture fine-grained road information for planning (Jia et al. 2025). Fig. 1 demonstrates the comparison between our method and current streaming-based learning methods(Yuan et al. 2024). To the best of our knowledge, achieving this objective presents two primary challenges: (1) Consistent positional embedding. Current streaming-based methods exhibit deficiencies in their positional embedding design for stream queries. Furthermore, certain lane topology reasoning approaches (Li et al. 2023a, b) suffer from inconsistency between reference points and positional embedding updates. (2) Temporal multiple attribute learning for lane segments. Topology reasoning between lanes is highly sensitive to the precise localization of lane segments (Fu et al. 2025a) and projection errors make it challenging to maintain consistent localization and category of lane segments across temporal propagation.
@@
To address the aforementioned challenges, we propose TopoStreamer, a novel temporal perception framework for lane segment topology reasoning. To strengthen positional embedding consistency, we augment the heads-to-regions mechanism (Li et al. 2023b) through dynamic explicit positional encoding across successive decoder layers. This design progressively injects updated positional information, thereby enhancing query updating with latest spatial learning. Furthermore, we introduce multiple streaming attribute constraints and a lane segment denoising module—both specifically optimized for lane segmentation tasks to reinforce temporal coherence and enable comprehensive feature representation learning. We also propose a new metric to evaluate lane boundary classification accuracy, a measure for autonomous vehicle lane-changing decision-making systems.
In summary, the primary contributions are outlined as follows:
  • We present TopoStreamer, a novel temporal lane segment perception method for lane topology reasoning in autonomous driving.
@@
  • Three novel modules have been proposed, including streaming attribute constraints for lane segments in temporal propagation, a dynamic lane boundary positional encoding module to enhance positional learning, and a lane segment denoising module to improve feature representation learning for position estimation, topological connections, and category prediction.
  • Extensive experiments conducted on lane segment benchmark OpenLane-V2 (Wang et al. 2024a) demonstrate SOTA performance of TopoStreamer in lane segment topology reasoning.
@@
Related Work
HD Map and Lane Topology Reasoning
Traditional high-definition (HD) map reconstruction primarily relies on SLAM-based methods (Zhang, Singh et al. 2014; Shan and Englot 2018), which incur substantial costs in manual annotation and map updates. With recent advancements in bird’s-eye view (BEV) perception and detection frameworks, offering improved efficiency and performance, the research focus has shifted towards vectorized HD map learning approaches. HDMapNet (Li et al. 2022b) generates HD semantic maps from multi-modal sensor data. However, extra post-processing is required to obtain vectorized representations. To generate vectorized map directly, VectorMapNet (Liu et al. 2023) predicts map elements as a set of polylines. The MapTR series (Liao et al. 2022, 2023) propose precise map element modeling and stabilizes learning via a hierarchical query-based anchor initialization mechanism. Unlike online HD map methods that primarily focus on drivable boundaries, our method concentrate on lane topology reasoning to perceive drivable trajectories (centerlines) and their topological relationships. STSU (Can et al. 2021) predicts an ordered lane graph to represent the traffic flow in the BEV. Subsequent research (Wu et al. 2023; Li et al. 2023a) has explored centerline topology using diverse model architectures on the OpenLane-V2 benchmark. To address endpoint misalignment issues in topology prediction, TopoLogic introduces dual constraints: distance-aware and similarity-aware optimization objectives. LaneSegNet (Li et al. 2023b) proposes lane segment perception to enhance the complete description of map. TopoPoint (Fu et al. 2025b) proposes Point-Lane interaction to learn accurate endpoints for reasoning. However, aforementioned methods overlook the potential benefits of temporal consistency for lane perception. In this work, we propose temporal-aware lane segment learning.
Temporal 3D Object Detection
In open-world scenarios, single-frame 3D detection faces challenges stemming from inaccurate pose estimation, occlusion, and adverse weather conditions. To overcome these limitations, recent advancements have incorporated long-term memory to store different feature. BEVFormerv2 (Yang et al. 2023) and BEVDet4D (Huang and Huang 2022) stack BEV features from historical frames. Sparse4D (Lin et al. 2022) and PETRv2 (Liu et al. 2022) design sparse fusion on images feature to avoid dense perspective transformation. StreamPETR (Wang et al. 2023) and Sparse4Dv3 (Lin et al. 2023) propagate historical information in query feature frame by frame. The temporal detection also shows impressive results in HD map learning. StreamMapNet (Yuan et al. 2024) proposes a streaming-based framework to warp and fuse the BEV features, and top-k reliable queries are selected to propagate. MapTracker (Chen et al. 2024) propose a tracking-based temporal fusion framework. It fuses the BEV and query features with distance strides to ensure extended-range consistency. SQD-MapNet proposes a denoising method for map elements to address the issue of information loss at the boundaries of BEV grid. Different from these methods, our distinctive contribution lies in the introduction of customized enhancements specifically designed for the more complex task of lane segment perception. This not only facilitates comprehensive road network understanding but also addresses the critical gap in positional embedding (PE) modeling within existing temporal map learning methods.
Query Denoising
Adding noise and performing denoising during training has been proven to accelerate the training process and enhance the model’s capabilities in both classification and regression. DN-DETR (Li et al. 2022a) introduces a denoising part apart from matching part as auxiliary supervision. DINO (Zhang et al. 2022) introduces a contrastive training to distinguish hard DN samples. SQD-MapNet (Wang et al. 2024b) proposes a stream query denoising to address the issue of map element truncation at boundaries caused by pose changes. To learn a comprehensive road network, we predict multiple attributes into a single lane segment query. To predict these attributes accurately, we design a tailored denoising learning strategy specifically for lane segments.
Method
@@
Given surrounding multi-view images $\mathbf{I}$, our goal is to predict 3D position, class attributes and topology of lane segments. Each lane segment is composed of a centerline $\mathbf{L}^{c}=(\mathbf{P},\mathit{C})$ , a left lane boundary $\mathbf{L}^{l}=(\mathbf{P},\mathit{T})$, and a right lane boundary $\mathbf{L}^{r}=(\mathbf{P},\mathit{T})$. $\mathbf{P}$ denotes an ordered set of points $\mathbf{P}=\{(x_{i},y_{i},z_{i})\}|^{M}_{i=1}$, where M is a preset number of points. In fact, we can obtain the boundary coordinates simply by predicting an offset and applying it to the centerline coordinates. $\mathit{C}$ indicates the lane segment class, which includes categories such as road lines and pedestrian crossings. $\mathit{T}$ denotes the boundary class, which can be dashed, solid, or non-visible. The connectivity topology is indicated by an adjacency matrix $\mathbf{A}$ (Can et al. 2021).
Overall Architecture
@@
The overall architecture of TopoStreamer is illustrated in Fig. 2. First, the surrounding multi-view images are processed by the PV-to-BEV encoder (Li et al. 2022c; He et al. 2016; Lin et al. 2017a) to generate BEV features $\textbf{F}_{bev}\in\mathbb{R}^{H\times W\times C}$, where C, H, W represent the number of feature channels, height, and width, respectively. These current BEV features are then fused with past BEV features. A DN query generator provides denoising (DN) queries $\textbf{Q}^{D}\in\mathbb{R}^{N\times C}$, DN center reference points $\mathbf{R^{D}_{C}}\in\mathbb{R}^{N\times M\times 3}$, and DN boundary reference points $\mathbf{R^{D}_{B}}\in\mathbb{R}^{N\times M\times 3}$. Next, a transformer-based decoder (Zhu et al. 2020) refines the DN, stream, and initialized queries $\{\textbf{Q}^{D},\textbf{Q}^{S},\textbf{Q}^{I}\}$. The BEV features, along with DN, stream, and initialized boundary reference points $\{\mathbf{R^{D}_{B}},\mathbf{R^{S}_{B}},\mathbf{R^{I}_{B}}\}$, are subsequently fed into the lane attention (Li et al. 2023b) along with the corresponding queries for further processing. A dynamic lane boundary positional encoding module encodes these boundary reference points into positional embeddings $\textbf{F}_{pe}\in\mathbb{R}^{N\times M\times C}$, injecting positional information into the queries layer by layer. Meanwhile, the DN, stream, and initialized center reference points $\{\mathbf{R^{D}_{C}},\mathbf{R^{S}_{C}},\mathbf{R^{I}_{C}}\}$ are utilized for prediction refinement (Zhu et al. 2020). The updated DN queries are used for lane segment denoising, while the stream and initialized queries are employed by the prediction heads to generate the lane graph. Additionally, past BEV features, queries and reference points are stored in a stream memory, enabling temporal propagation.
Temporal Propagation for Lane Segment
@@
Since lane segments remain stationary in geodetic coordinate while ego-vehicle poses change relative to them, we can utilize the detection results from the previous frame combined with the vehicle’s ego-motion to establish reference positions for subsequent frame predictions (Yuan et al. 2024). First, we warp the BEV features from the past frame to fuse with the BEV features from the current frame by a Gated Recurrent Unit (GRU) (Chung et al. 2014):
@@
Streaming Attribute Constraints. Conventional approaches (Yuan et al. 2024; Wang et al. 2024b; Chen et al. 2024) typically apply transformation loss to the converted coordinates to facilitate the learning of coordinate transformation. Since lane segments inherently possess multiple attributes, maintaining their temporal consistency requires more sophisticated constraints than simple coordinate transformation loss. To address this, we develop a comprehensive set of streaming attribute constraints. We employ MLPs to predict lane segment coordinate $\tilde{\mathbf{L}}_{t}=\text{Concat}(\tilde{\mathbf{L}}_{t}^{c},\tilde{\mathbf{L}}_{t}^{l},\tilde{\mathbf{L}}_{t}^{r})$ lane segment class $\tilde{\mathit{C}}_{t}$, boundary class $\tilde{\mathit{T}}_{t}$ and BEV semantic mask $\tilde{\mathbf{M}}_{t}$ from stream queries $\textbf{Q}_{t}^{S}$. Then, the streaming attribute constraints are represented as:
@@
where, for brevity, we omit the weights for each loss term. $\mathbf{L}_{t}$, $\mathit{T}_{t}$, $\mathit{C}_{t}$ and $\mathbf{M}_{t}$ are GT annotations transformed from T-1 frame to T frame.
Lossless Streaming Supervision. Existing approaches (Yuan et al. 2024; Wang et al. 2024b) utilize GT annotations from the past frame, transformed via pose estimation, to supervise the transformation loss for the subsequent frame. However, this method inevitably leads to information loss at BEV boundary regions as shown in (Wang et al. 2024b). To address this limitation, we track the unique IDs of positive instances matched through Hungarian assignment for stream queries, thereby providing lossless supervision. This is made possible by OpenLane-V2’s provision of unique instance IDs. For datasets that do not provide IDs, we can also use mask matching (Chen et al. 2024) to identify ID associations across frames.
Figure 3: Comparison of PE: (a) current streaming-based approaches (Yuan et al. 2024), (b) recent single-frame detection methods (Li et al. 2023b), and (c) our proposed method.
Dynamic Lane Boundary PE
As shown in Fig. 3, current temporal approaches (Yuan et al. 2024; Wang et al. 2024b; Chen et al. 2024) neglect the learning of positional embeddings, leading to inaccurate spatial localization. Furthermore, a critical updating inconsistency exists in recent single-frame detection methods (Li et al. 2023b, a). Because the initialized positional embeddings remain static, while the reference points are updated layer by layer. In the temporal propagation process, when some initial queries are substituted with stream queries, combining them with static PE could result in incompatible feature integration due to mismatches. The static PE refers to the PE that is not updated between layers during the forward pass. To address these problems, we enhance the heads-to-regions sampling module (Li et al. 2023b) by a dynamic lane boundary PE modeling. We apply point-wise positional encoding (Liu et al. 2024) to the boundary reference points to generate positional embeddings. We duplicate the queries to align with the number of boundary reference points. After the self-attention, we combine the positional embeddings with the corresponding queries. Subsequently, the queries interact with and BEV features through lane attention (LA) (Li et al. 2023b) with boundary reference points sampling. Finally, a MLP is used to merge the duplicated queries. This process can be represented as:

### v3

@@ Abstract
Lane segment topology reasoning constructs a comprehensive road network by capturing the topological relationships between lane segments and their semantic types. This enables end-to-end autonomous driving systems to perform road-dependent maneuvers such as turning and lane changing. However, the limitations in consistent positional embedding and temporal multiple attribute learning in existing methods hinder accurate road network reconstruction. To address these issues, we propose TopoStreamer, an end-to-end temporal perception model for lane segment topology reasoning. Specifically, TopoStreamer introduces three key improvements: streaming attribute constraints, dynamic lane boundary positional encoding, and lane segment denoising. The streaming attribute constraints enforce temporal consistency in both centerline and boundary coordinates, along with their classifications. Meanwhile, dynamic lane boundary positional encoding enhances the learning of up-to-date positional information within queries, while lane segment denoising helps capture diverse lane segment patterns, ultimately improving model performance. Additionally, we assess the accuracy of existing models using a lane boundary classification metric, which serves as a crucial measure for lane-changing scenarios in autonomous driving. On the OpenLane-V2 dataset, TopoStreamer demonstrates considerable improvements over state-of-the-art methods, achieving substantial performance gains of +3.0% mAP in lane segment perception and +1.7% OLS in centerline perception tasks. Code is accessible at https://github.com/YimingYang23/TopoStreamer.
1 Introduction
@@
Perception serves as a crucial component in end-to-end autonomous driving (Li et al., 2024b; Yang et al., 2025b), providing essential road priors for planning. Existing HD map learning and lane topology reasoning methods primarily focus on frame-by-frame detection (Li et al., 2023b; Liao et al., 2022). This approach fails to account for instance consistency across consecutive frames, making it susceptible to missed detections due to occlusions and high-speed movements (Yuan et al., 2024). Such limitations significantly hinder continuous and smooth decision-making and maneuvers. To comprehensively leverage temporal information, streaming-based methods (Yuan et al., 2024; Wang et al., 2024b; Wu et al., 2025) propose memory-based temporal propagation to establish long-term frame associations. Specifically, these approaches leverage the ego-vehicle pose to predict the probable positions of road instances in subsequent frames. However, these methods fail to capture sufficient road information for planning. This inspired us to introduce a temporal mechanism in lane topology reasoning, which we can leverage perception and topology reasoning results from previous frames to predict current frame outcomes and capture fine-grained road information for planning (Jia et al., 2025). Fig. 1 demonstrates the comparison between our method and current streaming-based learning methods(Yuan et al., 2024). To the best of our knowledge, achieving this objective presents two primary challenges: (1) Consistent positional embedding. Current streaming-based methods exhibit deficiencies in their positional embedding design for stream queries. Furthermore, certain lane topology reasoning approaches (Li et al., 2023a; b) suffer from inconsistency between reference points and positional embedding updates. (2) Temporal multiple attribute learning for lane segments. Topology reasoning between lanes is highly sensitive to the precise localization of lane segments (Fu et al., 2025a) and projection errors make it challenging to maintain consistent localization and category of lane segments across temporal propagation.
Figure 1: Comparsion between current streaming-based map learning methods (Yuan et al., 2024) and our TopoStreamer. TopoStreamer delivers more fine-grained road information through streaming perception of lane segments, which is vital for planning.
@@
To address the aforementioned challenges, we propose TopoStreamer, a novel temporal perception framework for lane segment topology reasoning. To strengthen positional embedding consistency, we augment the heads-to-regions mechanism (Li et al., 2023b) through dynamic explicit positional encoding across successive decoder layers. This design progressively injects updated positional information to enhance query updating with latest spatial learning. Furthermore, we introduce multiple streaming attribute constraints and a lane segment denoising module to reinforce temporal coherence and enable the learning of diverse patterns in lane segments. We also propose a new metric to evaluate the lane boundary classification accuracy, a measure for autonomous vehicle lane-changing decision-making systems.
@@
Contributions: (1) We present TopoStreamer, a novel temporal lane segment perception method for lane topology reasoning in autonomous driving. (2) Three novel modules have been proposed, including streaming attribute constraints for lane segments in temporal propagation, a dynamic lane boundary positional encoding module to enhance positional learning, and a lane segment denoising module for the learning of diverse patterns in lane segments. (3) Extensive experiments conducted on lane segment benchmark OpenLane-V2 (Wang et al., 2024a) demonstrate SOTA performance of TopoStreamer in lane topology reasoning.
2 Related Work
2.1 HD Map and Lane Topology Reasoning
Traditional high-definition (HD) map reconstruction primarily relies on SLAM-based methods (Zhang et al., 2014; Shan & Englot, 2018), which incur substantial costs in manual annotation and map updates. With recent advancements in bird’s-eye view (BEV) perception and detection frameworks, offering improved efficiency and performance, the research focus has shifted towards vectorized HD map learning approaches. HDMapNet (Li et al., 2022b) generates HD semantic maps from multi-modal sensor data. However, extra post-processing is required to obtain vectorized representations. To generate vectorized map directly, VectorMapNet (Liu et al., 2023) predicts map elements as a set of polylines. The MapTR series (Liao et al., 2022; 2023) propose precise map element modeling and stabilizes learning via a hierarchical query-based anchor initialization mechanism. Unlike online HD map methods that primarily focus on drivable boundaries, our method concentrate on lane topology reasoning to perceive drivable trajectories (centerlines) and their topological relationships. STSU (Can et al., 2021) predicts an ordered lane graph to represent the traffic flow in the BEV. Subsequent research (Wu et al., 2023; Li et al., 2023a) has explored centerline topology using diverse model architectures on the OpenLane-V2 benchmark. To address endpoint misalignment issues in topology prediction, TopoLogic introduces dual constraints: distance-aware and similarity-aware optimization objectives. LaneSegNet (Li et al., 2023b) proposes lane segment perception to enhance the complete description of map. TopoPoint (Fu et al., 2025b) proposes Point-Lane interaction to learn accurate endpoints for reasoning. However, aforementioned methods overlook the potential benefits of temporal consistency for lane perception. In this work, we propose temporal-aware lane segment learning.
2.2 Temporal 3D Object Detection
In open-world scenarios, single-frame 3D detection faces challenges stemming from inaccurate pose estimation, occlusion, and adverse weather conditions. To overcome these limitations, recent advancements have incorporated long-term memory to store different feature. BEVFormerv2 (Yang et al., 2023) and BEVDet4D (Huang & Huang, 2022) stack BEV features from historical frames. Sparse4D (Lin et al., 2022) and PETRv2 (Liu et al., 2022) design sparse fusion on images feature to avoid dense perspective transformation. StreamPETR (Wang et al., 2023) and Sparse4Dv3 (Lin et al., 2023) propagate historical information in query feature frame by frame. The temporal detection also shows impressive results in HD map learning. StreamMapNet (Yuan et al., 2024) proposes a streaming-based framework to warp and fuse the BEV features, and top-k reliable queries are selected to propagate. MapTracker (Chen et al., 2024) propose a tracking-based temporal fusion framework. It fuses the BEV and query features with distance strides to ensure extended-range consistency. SQD-MapNet proposes a denoising method for map elements to address the issue of information loss at the boundaries of BEV grid. Different from these methods, our distinctive contribution lies in the introduction of customized enhancements specifically designed for the more complex task of lane segment perception. This not only facilitates comprehensive road network understanding but also addresses the critical gap in positional embedding (PE) modeling within existing temporal map learning methods.
2.3 Query Denoising
Adding noise and performing denoising during training has been proven to accelerate the training process and enhance the model’s capabilities in both classification and regression. DN-DETR (Li et al., 2022a) introduces a denoising part apart from matching part as auxiliary supervision. DINO (Zhang et al., 2022) introduces a contrastive training to distinguish hard DN samples. SQD-MapNet (Wang et al., 2024b) proposes a stream query denoising to address the issue of map element truncation at boundaries caused by pose changes. To learn a comprehensive road network, we predict multiple attributes into a single lane segment query. To predict these attributes accurately, we design a tailored denoising learning strategy specifically for lane segments.
@@ 2.3 Query Denoising
3 Method
@@
Given surrounding multi-view images $\mathbf{I}$, our goal is to predict 3D position, class attributes and topology of lane segments. Each lane segment is composed of a centerline $\mathbf{L}^{c}=(\mathbf{P},\mathit{C})$ , a left lane boundary $\mathbf{L}^{l}=(\mathbf{P},\mathit{T})$, and a right lane boundary $\mathbf{L}^{r}=(\mathbf{P},\mathit{T})$. $\mathbf{P}$ denotes an ordered set of points $\mathbf{P}=\{(x_{i},y_{i},z_{i})\}|^{M}_{i=1}$, where $M$ is a preset number of points. In fact, we can obtain the boundary coordinates simply by predicting an offset and applying it to the centerline coordinates. $\mathit{C}$ indicates the lane segment class, which includes categories such as road lines and pedestrian crossings. $\mathit{T}$ denotes the boundary class, which can be dashed, solid, or non-visible. The connectivity topology is indicated by an adjacency matrix $\mathbf{A}$ (Can et al., 2021).
3.1 Overall Architecture
@@
The overall architecture of TopoStreamer is illustrated in Fig. 2. First, the surrounding multi-view images are processed by the PV-to-BEV encoder (Li et al., 2022c; He et al., 2016; Lin et al., 2017a) to generate BEV features $\textbf{F}_{bev}\in\mathbb{R}^{H\times W\times C}$, where C, H, W represent the number of feature channels, height, and width, respectively. These current BEV features are then fused with past BEV features. A DN query generator provides denoising (DN) queries $\textbf{Q}^{D}\in\mathbb{R}^{N\times C}$, DN center reference points $\mathbf{R^{D}_{C}}\in\mathbb{R}^{N\times M\times 3}$, and DN boundary reference points $\mathbf{R^{D}_{B}}\in\mathbb{R}^{N\times M\times 3}$. Next, a transformer-based decoder (Zhu et al., 2020) refines the DN, stream, and initialized queries $\{\textbf{Q}^{D},\textbf{Q}^{S},\textbf{Q}^{I}\}$. The BEV features, along with DN, stream, and initialized boundary reference points $\{\mathbf{R^{D}_{B}},\mathbf{R^{S}_{B}},\mathbf{R^{I}_{B}}\}$, are subsequently fed into the lane attention (Li et al., 2023b) along with the corresponding queries for further processing. A dynamic lane boundary positional encoding module encodes these boundary reference points into positional embeddings $\textbf{F}_{pe}\in\mathbb{R}^{N\times M\times C}$, injecting positional information into the queries layer by layer. Meanwhile, the DN, stream, and initialized center reference points $\{\mathbf{R^{D}_{C}},\mathbf{R^{S}_{C}},\mathbf{R^{I}_{C}}\}$ are utilized for prediction refinement (Zhu et al., 2020). The updated DN queries are used for lane segment denoising, while the stream and initialized queries are employed by the prediction heads to generate the lane graph. Additionally, past BEV features, queries and reference points are stored in a stream memory, enabling temporal propagation.
3.2 Temporal Propagation for Lane Segment
@@
Since lane segments remain stationary in geodetic coordinate while ego-vehicle poses change relative to them, we can utilize the detection results from the previous frame combined with the vehicle’s ego-motion to establish reference positions for subsequent frame predictions (Yuan et al., 2024). First, we warp the BEV features from the past frame to fuse with the BEV features from the current frame by a Gated Recurrent Unit (GRU) (Chung et al., 2014):
@@
Streaming Attribute Constraints. Conventional approaches (Yuan et al., 2024; Wang et al., 2024b; Chen et al., 2024) typically apply transformation loss to the converted coordinates to facilitate the learning of coordinate transformation. Since lane segments inherently possess multiple attributes, maintaining their temporal consistency requires more sophisticated constraints than simple coordinate transformation loss. To address this, we develop a comprehensive set of streaming attribute constraints. We employ MLPs to predict lane segment coordinate $\tilde{\mathbf{L}}_{t}=\text{Concat}(\tilde{\mathbf{L}}_{t}^{c},\tilde{\mathbf{L}}_{t}^{l},\tilde{\mathbf{L}}_{t}^{r})$ lane segment class $\tilde{\mathit{C}}_{t}$, boundary class $\tilde{\mathit{T}}_{t}$ and BEV semantic mask $\tilde{\mathbf{M}}_{t}$ from stream queries $\textbf{Q}_{t}^{S}$. Then, the streaming attribute constraints are represented as:
@@
where, for brevity, we omit the weights for each loss term. $\mathbf{L}_{t}$, $\mathit{T}_{t}$, $\mathit{C}_{t}$ and $\mathbf{M}_{t}$ are GT annotations transformed from T-1 frame to T frame. More details can be found in appendix.
Lossless Streaming Supervision. Existing approaches (Yuan et al., 2024; Wang et al., 2024b) utilize GT annotations from the past frame, transformed via pose estimation, to supervise the transformation loss for the subsequent frame. However, this method inevitably leads to information loss at BEV boundary regions as shown in (Wang et al., 2024b). To address this limitation, we track the unique IDs of positive instances matched through Hungarian assignment for stream queries, thereby providing lossless supervision. This is made possible by OpenLane-V2’s provision of unique instance IDs. For datasets that do not provide IDs, we can also use mask matching (Chen et al., 2024) to identify ID associations across frames.
3.3 Dynamic Lane Boundary PE
Figure 3: Comparison of PE: (a) current streaming-based approaches (Yuan et al., 2024), (b) recent single-frame detection methods (Li et al., 2023b), and (c) our proposed method.
As shown in Fig. 3, current temporal approaches (Yuan et al., 2024; Wang et al., 2024b; Chen et al., 2024) neglect the learning of positional embeddings, leading to inaccurate spatial localization. Furthermore, a critical updating inconsistency exists in recent single-frame detection methods (Li et al., 2023b; a). Because the initialized positional embeddings remain static, while the reference points are updated layer by layer. In the temporal propagation process, when some initial queries are substituted with stream queries, combining them with static PE could result in incompatible feature integration due to mismatches. The static PE refers to the PE that is not updated between layers during the forward pass. To address these problems, we enhance the heads-to-regions sampling module (Li et al., 2023b) by a dynamic lane boundary PE modeling. We apply point-wise positional encoding (Liu et al., 2024) to the boundary reference points to generate positional embeddings. We duplicate the queries to align with the number of boundary reference points. After the self-attention, we combine the positional embeddings with the corresponding queries. Subsequently, the queries interact with and BEV features through lane attention (LA) (Li et al., 2023b) with boundary reference points sampling. Finally, a MLP is used to merge the duplicated queries. This process can be represented as: